# PSformer: Parameter-efficient Transformer with Segment Attention for Time Series Forecasting

### v1

@@ 4 Experiment
Metric PSformer SAMformer TSMixer PatchTST MOMENT ModernTCN FEDformer GPT4TS Autoformer RLinear iTransformer MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 0.397 0.418 0.410 0.424 0.420 0.431 0.468 0.455 0.418 0.436 0.421 0.432 0.428 0.454 0.428 0.426 0.473 0.477 0.446 0.434 0.454 0.448 ETTh2 0.338 0.390 0.344 0.391 0.354 0.400 0.387 0.407 0.352 0.394 0.343 0.393 0.388 0.434 0.355 0.395 0.422 0.443 0.374 0.398 0.383 0.407 ETTm1 0.342 0.372 0.373 0.388 0.378 0.392 0.387 0.400 0.344 0.379 0.361 0.384 0.382 0.422 0.351 0.383 0.515 0.493 0.414 0.407 0.407 0.410 ETTm2 0.251 0.313 0.269 0.327 0.283 0.339 0.281 0.326 0.259 0.318 0.262 0.322 0.292 0.343 0.267 0.326 0.310 0.357 0.286 0.327 0.288 0.332 Weather 0.225 0.264 0.261 0.293 0.255 0.289 0.259 0.281 0.228 0.270 0.237 0.274 0.310 0.357 0.237 0.271 0.335 0.379 0.272 0.291 0.258 0.278 Electricity 0.162 0.255 0.181 0.275 0.198 0.296 0.205 0.290 0.165 0.260 0.160 0.255 0.207 0.321 0.167 0.263 0.214 0.327 0.219 0.298 0.178 0.270 Exchange 0.417 0.432 0.445 0.470 0.532 0.523 0.367 0.404 0.437 0.446 0.555 0.536 0.478 0.477 0.371 0.409 0.613 0.539 0.378 0.417 0.360 0.350 Traffic 0.400 0.274 0.425 0.297 0.439 0.315 0.481 0.304 0.415 0.293 0.414 0.283 0.604 0.372 0.414 0.295 0.617 0.384 0.626 0.378 0.428 0.282 Count 6 7 0 0 0 0 1 0 0 0 1 1 0 0 0 0 0 0 0 0 1 1
@@ 4.1 Results and Analysis
Summary of Experimental Results The main experimental results are reported in Table 2. PSformer achieved the best performance on 6 out of 8 major datasets in long-term time series forecasting tasks, demonstrating its strong predictive capabilities across various time series prediction problems. This success is attributed to its segment attention mechanism, which enhances the extraction of spatial-temporal information, and the parameter-sharing structure, which improves the model’s robustness. The complete experimental results are in Appendix B.2.
@@ 4.2 Ablation studies
The Effect of segments Numbers Since PSformer employs a non-overlapping patching to construct segments, the model’s performance is affected by the number of segments. Therefore, we tested the model’s performance with different segment numbers on two datasets, ETTh1 and ETTm1. Given that the input sequence length is fixed at 512, the number of segments must be a divisor of the sequence length. Consequently, we set the number of segments to 8, 16, 32, and 64, and test the model on four different forecasting horizons. The test results are shown in Table 3, which indicate that the number of segments impacts the model’s prediction accuracy to some extent. Across both datasets, a moderate number of segments (such as 32) tends to perform the best, particularly for both short and long forecasting horizons. For shorter horizons (96, 192), using 16 or 32 segments generally yields the lowest MSE, while fewer segments (8) often result in poorer performance. As the horizon increases (336, 720), 32 segments consistently lead to optimal results, indicating that this number balances the extraction of both local and global temporal features effectively.
@@
Num. of segments ETTh1 Num. of segments ETTm1 96 192 336 720 96 192 336 720 8 0.362 0.406 0.670 0.451 8 0.290 0.325 0.357 0.412 16 0.352 0.417 0.424 0.446 16 0.285 0.323 0.353 0.412 32 0.352 0.386 0.410 0.440 32 0.282 0.321 0.352 0.413 64 0.354 0.389 0.412 0.446 64 0.288 0.325 0.355 0.417
@@
Num. of Encoders ETTh1 Num. of Encoders ETTm1 96 192 336 720 96 192 336 720 1 0.352 0.385 0.411 0.440 1 0.288 0.324 0.356 0.414 2 0.355 0.392 0.418 0.443 2 0.284 0.323 0.356 0.415 3 0.355 0.391 0.416 0.443 3 0.282 0.321 0.352 0.413 4 0.355 0.389 0.416 0.440 4 0.282 0.321 0.353 0.416
Ablation of Parameter Sharing methods We investigate the impact of parameter-sharing mechanism on the model performance. In addition to the default parameter-sharing approach, which shares parameters only within encoder (In-Encoder), we also test the following approaches: a. no parameter sharing, i.e., None; b. sharing parameters only across encoders (with different parameters used for the PS blocks within each encoder), i.e., Cross-Encoders; and c. sharing parameters both within and across encoders, i.e., ALL. We conducted experiments on the ETTm1 and Weather datasets, and the results are shown in Table 5. As can be seen, the In-Encoder method performs the best, followed by ALL, while None shows the worst performance. This indicates that the parameter-sharing mechanism contributes to improving the model performance.
@@
Sharing Methods Cross-Encoders In-Encoder ALL None
ETTm1 96 0.297 0.282 0.288 0.295
196 0.329 0.321 0.326 0.338
336 0.360 0.352 0.358 0.372
720 0.420 0.413 0.414 0.425
Weather 96 0.158 0.149 0.151 0.154
196 0.201 0.193 0.196 0.198
336 0.252 0.245 0.245 0.245
720 0.319 0.314 0.316 0.319
@@ 5 Conclusion and future work
In this work, we proposed the PSformer model for multivariate time series forecasting, which leverages segment attention technique to facilitate information transfer among time series variables and capture spatial-temporal dependencies. By employing parameter sharing, the model effectively improves parameter efficiency and reduces the risk of overfitting when the data size is relatively limited. This network structure, which combines parameter sharing with segment attention mechanism, achieves state-of-the-art performance on long-term multivariate forecasting tasks by enhancing model parameter efficiency and improving the utilization of channel-wise information. Future work can focus on applying this parameter sharing and segment attention technique to the development of pre-trained large models for time series forecasting, to overcome the issue of excessively large parameter counts in existing pre-trained models, and improve the capability of extracting information from multivariate time series. Additionally, exploring model architecture designs with parameter sharing to improve prediction performance still holds potential value.
Acknowledgments
We would like to thank Haixu Wu for his valuable feedback and suggestions for improving our manuscript prior to submission. His contributions were greatly appreciated and helped shape the final version of our work.
@@ Appendix A Experimental Configuration
A.1 Details of Baseline Settings
@@ A.2 Settings for PSformer
The learning rate adjustment strategy (lradj) is set to ”constant” for all experiments, with a fixed learning rate (LR) of $10^{-4}$. The loss function used in all experiments is Mean Squared Error (MSE). The batch size is set to 16 for most tasks, except for the Traffic dataset, where the batch size is 8. Each experiment runs for 300 epochs, with a patience value of 30 for early stopping. A fixed random seed of 1 is applied across all experiments to ensure reproducibility.
@@
Task-Dataset Encoder Num. Input Length $T$ Segment Num. $N$ lradj LR* Loss Batch Size Epochs patient random seed ETTh1 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTh2 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTm1 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTm2 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 Weather 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 Electricity 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 Traffic 3 512 32 constant $10^{-4}$ MSE 8 300 30 1 Exchange 1 512 32 constant $10^{-4}$ MSE 16 300 30 1
A.3 Model Size Comparison
@@
A.4 Running time Comparison
@@
A.5 Discuss about Positional Encoding
The reasons of eliminating positional encoding Similarly, SAMformer does not use positional encoding because it applies attention to the channel dimension, where there is no strict sequential relationship between channels. Although SegAtt is also a cross-channel structure, it involves local time series constructed by patches. We consider such local sequences as local representations (or tokens) rather than short sequences. The experimental results also demonstrate that this structure can similarly reduce the dependency on positional encoding while achieving good performance.
A.6 Sharpness-Aware Minimization (SAM)
@@
Appendix B More Results and Analysis
B.1 Investigation of hyper-parameter $\rho$
The Effect of $\rho$ Since we employed SAM to ensure training stability, we also conducted sensitivity tests on the hyper-parameter $\rho$ in SAM. We divided the range of $\rho$ from 0 to 1 into 10 equal parts and tested its effect on model prediction performance across the ETTh1, ETTm1, ETTm2, and Weather datasets. The results are shown in Figure 3. It can be observed that as the parameter $\rho$ gradually increases in SAM can improve the model’s prediction performance to some extent. However, if $\rho$ is set too large, it may degrade the model’s performance. When selecting $\rho$, it’s important to consider the dataset’s complexity and noise levels, as well as the model’s architecture. For more complex datasets or larger models, a slightly larger $\rho$ can help smooth the loss landscape and improve generalization. Further, $\rho$ should also be balanced with the learning rate to avoid instability or performance degradation. As a comparison, we also report the $\rho^{*}$ used by PSformer, SAMformer and TSMixer in the Table 10.

### v2

@@ 3.2 Model Structure
Metric PSformer SAMformer TSMixer PatchTST MOMENT ModernTCN FEDformer GPT4TS Autoformer RLinear iTransformer MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE ETTh1 0.397 0.418 0.410 0.424 0.420 0.431 0.468 0.455 0.418 0.436 0.421 0.432 0.428 0.454 0.428 0.426 0.473 0.477 0.446 0.434 0.454 0.448 ETTh2 0.338 0.390 0.344 0.391 0.354 0.400 0.387 0.407 0.352 0.394 0.343 0.393 0.388 0.434 0.355 0.395 0.422 0.443 0.374 0.398 0.383 0.407 ETTm1 0.342 0.372 0.373 0.388 0.378 0.392 0.387 0.400 0.344 0.379 0.361 0.384 0.382 0.422 0.351 0.383 0.515 0.493 0.414 0.407 0.407 0.410 ETTm2 0.251 0.313 0.269 0.327 0.283 0.339 0.281 0.326 0.259 0.318 0.262 0.322 0.292 0.343 0.267 0.326 0.310 0.357 0.286 0.327 0.288 0.332 Weather 0.225 0.264 0.261 0.293 0.255 0.289 0.259 0.281 0.228 0.270 0.237 0.274 0.310 0.357 0.237 0.271 0.335 0.379 0.272 0.291 0.258 0.278 Electricity 0.162 0.255 0.181 0.275 0.198 0.296 0.205 0.290 0.165 0.260 0.160 0.255 0.207 0.321 0.167 0.263 0.214 0.327 0.219 0.298 0.178 0.270 Exchange 0.358 0.399 0.445 0.470 0.532 0.523 0.367 0.404 0.437 0.446 0.555 0.536 0.478 0.477 0.371 0.409 0.613 0.539 0.378 0.417 0.360 0.403 Traffic 0.400 0.274 0.425 0.297 0.439 0.315 0.481 0.304 0.415 0.293 0.414 0.283 0.604 0.372 0.414 0.295 0.617 0.384 0.626 0.378 0.428 0.282 Count 7 8 0 0 0 0 1 0 0 0 1 1 0 0 0 0 0 0 0 0 1 1
4 Experiment
Datasets In this paper, we focus on the long-term time series forecasting. We follow the time series forecasting work in (Ilbert et al., 2024) and use 8 mainstream datasets to evaluate the performance of our proposed PSformer model. As shown in Table 1, these datasets include 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), as well as Weather, Traffic, Electricity, and Exchange. These datasets have been used as benchmark evaluations in many previous time series forecasting studies.
Baselines We select state-of-the-art (SOTA) models in the field of long-term time series forecasting, including not only Transformer-based models but also large models and other SOTA models. Specifically, baselines include (1) Transformer-based model: SAMformer (Ilbert et al., 2024), iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), (2) Pretrained Large model: MOMENT (Goswami et al., 2024), GPT4TS (Zhou et al., 2023), (3) TCN-based model: ModernTCN (donghao & wang xue, 2024) and (4) MLP-based methods: TSMixer (Chen et al., 2023b), RLinear (Li et al., 2023). Additionally, we provide more baselines for a comprehensive comparison, including TimeMixer (Wang et al., 2024a), CrossGNN (Huang et al., 2023), MICN (Wang et al., 2023), TimesNet (Wu et al., 2023), FITS (Xu et al., 2024), Crossformer (Zhang & Yan, 2023), PDF (Dai et al., 2024), and TimeLLM (Jin et al., 2024) Further details about these baselines can be found in Section B.8.
Experimental Settings The input time series length $T$ is set to 512, and four different prediction lengths $H\in\{96,192,336,720\}$ are used. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE). We train our constructed models using the SAM optimization technique as in (Ilbert et al., 2024). Setting the look-back window of RevIN to 16 for the Exchange dataset, more details about this setting can be found in Section B.7. We collect baseline results of SAMformer, TSMixer, FEDformer and Autoformer from (Ilbert et al., 2024), in which SAMformer and TSMixer are also trained with SAM, and the results of iTransformer, RLinear and PatchTST from (Liu et al., 2024b). We test ModernTCN with the fixed input length of 512 following (donghao & wang xue, 2024). For the pre-trained large models MOMENT and GPT4TS, the results are collected from (Goswami et al., 2024). More details about the baseline settings can be found in Section A.2. Besides, We also provide results under uni-variate series and unrelated variate in the Section B.6.
@@ 4.1 Results and Analysis
Summary of Experimental Results The main experimental results are reported in Table 2. PSformer achieved the best performance on 7 out of 8 major datasets in long-term time series forecasting tasks, demonstrating its strong predictive capabilities across various time series prediction problems. This success is attributed to its segment attention mechanism, which enhances the extraction of spatial-temporal information, and the parameter-sharing structure, which improves the model’s robustness. The complete experimental results are in Appendix B.2. We also provide additional visualization results to analyze the attention maps across spatial and temporal dimensions in the Section B.5.
@@ 4.2 Ablation studies
The Effect of segments Numbers Since PSformer employs a non-overlapping patching to construct segments, the model’s performance is affected by the number of segments. Therefore, we tested the model’s performance with different segment numbers on two datasets, ETTh1 and ETTm1. Given that the input sequence length is fixed at 512, the number of segments must be a divisor of the sequence length. Consequently, we set the number of segments to 8, 16, 32, and 64, and test the model on four different forecasting horizons. The test results are shown in Table 3, which indicate that the number of segments impacts the model’s prediction accuracy to some extent. Across both datasets, a moderate number of segments (such as 32) tends to perform the best, particularly for both short and long forecasting horizons. For shorter horizons (96, 192), using 16 or 32 segments generally yields the lowest MSE, while fewer segments (8) often result in poorer performance. As the horizon increases (336, 720), 32 segments consistently lead to optimal results, indicating that this number balances the extraction of both local and global temporal features effectively. Besides, we provide results and analysis about the impact of the segment number across more datasets in the Section B.12.
@@
# of Seg. ETTh1 ETTm1 96 192 336 720 96 192 336 720 8 0.362 0.406 0.670 0.451 0.290 0.325 0.357 0.412 16 0.352 0.417 0.424 0.446 0.285 0.323 0.353 0.412 32 0.352 0.386 0.410 0.440 0.282 0.321 0.352 0.413 64 0.354 0.389 0.412 0.446 0.288 0.325 0.355 0.417
# of Encoder ETTh1 ETTm1 96 192 336 720 96 192 336 720 1 0.352 0.385 0.411 0.440 0.288 0.324 0.356 0.414 2 0.355 0.392 0.418 0.443 0.284 0.323 0.356 0.415 3 0.355 0.391 0.416 0.443 0.282 0.321 0.352 0.413 4 0.355 0.389 0.416 0.440 0.282 0.321 0.353 0.416
@@
Ablation of Parameter Sharing methods We investigate the impact of parameter-sharing mechanism on the model performance. In addition to the default parameter-sharing approach, which shares parameters only within encoder (In-Encoder), we also test the following approaches: a. no parameter sharing, i.e., None; b. sharing parameters only across encoders (with different parameters used for the PS blocks within each encoder), i.e., Cross-Encoders; and c. sharing parameters both within and across encoders, i.e., ALL. We conducted experiments on the ETTm1 and Weather datasets, and the results are shown in Table 5. As can be seen, the In-Encoder method performs the best, followed by ALL, while None shows the worst performance. This indicates that the parameter-sharing mechanism contributes to improving the model performance. Furthermore, we provide a comparison of convergence rates between parameter sharing and non-parameter sharing in Section B.9, as well as a comparison of parameter savings achieved by parameter sharing in Section B.10 and Section A.4.
@@
Sharing Methods Cross-Encoders In-Encoder ALL None ETTm1 96 0.297 0.282 0.288 0.295 196 0.329 0.321 0.326 0.338 336 0.360 0.352 0.358 0.372 720 0.420 0.413 0.414 0.425 Weather 96 0.158 0.149 0.151 0.154 196 0.201 0.193 0.196 0.198 336 0.252 0.245 0.245 0.245 720 0.319 0.314 0.316 0.319
@@ 5 Conclusion and Future Work
In this work, we proposed the PSformer model for multivariate time series forecasting, which leverages segment attention technique to facilitate information transfer among time series variables and capture spatial-temporal dependencies. By employing parameter sharing, the model effectively improves parameter efficiency and reduces the risk of overfitting when the data size is relatively limited. Overall, this designed network structure achieves state-of-the-art performance on long-term multivariate forecasting tasks by enhancing model parameter efficiency and improving the utilization of channel-wise information. Future work can focus on applying this parameter sharing and segment attention technique to the development of pre-trained large models for time series forecasting, to overcome the issue of excessively large parameter counts in existing pre-trained models, and improve the capability of extracting information from multivariate time series.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experimental Configuration
A.1 Hardware
All experiments were conducted on two servers, each equipped with an 80GB NVIDIA A100 GPU and 4 Intel Xeon Gold 5218 CPUs.
A.2 Details of Baseline Settings
@@ A.3 Settings for PSformer
The learning rate adjustment strategy (schedule) is set to “constant” for all experiments, with a fixed learning rate (LR) of $10^{-4}$. The loss function used in all experiments is Mean Squared Error (MSE). The batch size is set to 16 for most tasks, except for the Traffic dataset, where the batch size is 8. Each experiment runs for 300 epochs, with a patience value of 30 for early stopping. A fixed random seed of 1 is applied across all experiments to ensure reproducibility.
@@
Task-Dataset Encoder Num. Input Length $T$ Segment Num. $N$ schedule LR* Loss Batch Size Epochs patient random seed ETTh1 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTh2 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTm1 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 ETTm2 1 512 32 constant $10^{-4}$ MSE 16 300 30 1 Weather 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 Electricity 3 512 32 constant $10^{-4}$ MSE 16 300 30 1 Traffic 3 512 32 constant $10^{-4}$ MSE 8 300 30 1 Exchange 1 512 32 constant $10^{-4}$ MSE 16 300 30 1
A.4 Model Size Comparison
@@
A.5 Running time Comparison
@@
A.6 Discuss about Positional Encoding
The reasons of eliminating positional encoding On the one hand, the SAMformer does not use positional encoding because it applies attention to the channel dimension, where there is no strict sequential relationship between channels. Although SegAtt in PSformer is also a cross-channel structure, it involves local time series constructed by patches. We consider such local sequences as local representations (or tokens) rather than short sequences. The experimental results also demonstrate that this structure can similarly reduce the dependency on positional encoding while achieving good performance.
On the other hand, we conducted a set of comparative experiments using positional encoding to better illustrate its impact. Specifically, we tested the encoding performance under two different data transformation modes: pos emb (time series), where positional encoding is applied to the original time series before dimension transformation; and pos emb (segment), where positional encoding is applied to the transformed segments. The default case, No pos emb, refers to the absence of positional encoding.
To highlight the differences between seasonal vs. non-seasonal and stationary vs. non-stationary characteristics, we selected the ETTh1 and ETTh2 datasets (relatively seasonal and stable) as well as the Exchange dataset (relatively non-seasonal and non-stationary). The experimental results are shown in the Table 10.
The degraded performance of pos emb (time series) might be due to the incompatibility of the positional encoding with dimension transformation, as the original temporal order is lost in the segment dimension, making it unsuitable for dot-product attention calculations. On the other hand, pos emb (segment) shows smaller changes compared to the No pos emb case, but the performance still deteriorates slightly. This suggests that the significance of positional encoding in the context of multivariate time series forecasting might need to be re-evaluated, as there are fundamental differences between NLP and time series data when applying attention mechanisms.
Table 10: Performance comparison with different positional embedding methods.
Dataset Horizon Pos emb (time series) Pos emb (segment) No pos emb ETTh1 96 0.378 0.353 0.352 192 0.412 0.388 0.385 ETTh2 96 0.298 0.276 0.272 192 0.352 0.338 0.335 Exchange Rate 96 0.189 0.095 0.091 192 0.314 0.201 0.197 336 0.525 0.370 0.345 720 1.574 1.041 1.036
A.7 Sharpness-Aware Minimization (SAM)
@@
Appendix B More Results and Analysis
B.1 Investigation of hyper-parameter $\rho$
The Effect of $\rho$. Since we employed SAM to ensure training stability, we also conducted sensitivity tests on the hyper-parameter $\rho$ in SAM. We divided the range of $\rho$ from 0 to 1 into 10 equal parts and tested its effect on model prediction performance across the ETTh1, ETTm1, ETTm2, and Weather datasets. The results are shown in Figure 3. It can be observed that as the parameter $\rho$ gradually increases in SAM can improve the model’s prediction performance to some extent. However, if $\rho$ is set too large, it may degrade the model’s performance. When selecting $\rho$, it’s important to consider the dataset’s complexity and noise levels, as well as the model’s architecture. For more complex datasets or larger models, a slightly larger $\rho$ can help smooth the loss landscape and improve generalization. Further, $\rho$ should also be balanced with the learning rate to avoid instability or performance degradation. As a comparison, we also report the $\rho^{*}$ used by PSformer, SAMformer and TSMixer in the Table 11.